Experimental Summary

There is now published or preliminary data on

• Top Cross Section

σt t = 7.6{+2.4 \atop −2.0} pb (CDF )
σt t = 6.3 ± 2.2 pb (D0 / )

• Top Mass

Mtop = 176 ± 8 ± 10 GeV/c2 (CDF )
Mtop = 199{+19 \atop −21} ± 22 GeV/c2 (D0 / )

• Top branching fractions BR(t → W + b)

R = 0.87{+0.13 \atop −0.30} (stat ){+0.13 \atop −0.11} (syst )

Much remains to be done....a not-so-complete list

• Search for rare decays (eg. t → Z/γq)

• Study detailed properties of the t quark – Production properties (eg, σt t, ALR)

• Search for new gauge bosons coupling to tb̄ – SUSY Higgs

– Longitudinal W ’s

• Look for new physics coupling to t or t t

These will require high statistics samples of t t events

How To Optimise Top Physics?

More statistics an obvious place to push

• Increase integrated luminosity ∫ L dt – More running time

– Upgraded Tevatron

– LHC in next decade

• Improved detector tagging capability – b tagging via SVX works extremely well

– Increase coverage at CDF (only ∼ 50% now)

– Add capability to D0 /

• Increase collision energy – Increase from 1.8 to 2.0 TeV yields 40% more t t – LHC energy of 14 TeV yields ×100

Basic detection techniques hard to improve upon

• Sorting out jets remains messy

• With higher statistics, can choose cleaner subsamples – High pT top quark events

– Allows complementary analysis techniques

Future Hadron Machines

Have to consider upgraded Tevatron and LHC

Machine  Year Energy  ∫ L dt
(TeV)  (fb−1)
Tevatron Run I  1996  1.8  ∼ 0.15
Tevatron Run II  1999  2.0  1
Tevatron Run III  2003  2.0?  30 − 100?
LHC  2005  8-14  100 − 1000?

Use as framework to look at impact on top physics

• No imponderable technical hurdles to overcome

• Tevatron Run II is happening – Run III depends on running schedule and LHC

– Other proposals being kicked around

Tevatron Collider Development

The immediate next step is the Main Injector Project

• Improvements are Main Injector and p̄ source – Overcomes proton density limitation in Main Ring – Higher acceptance into Collider

• Typical luminosity will increase ×4 to 1032 cm−2s−1 – Bunch spacing goes from 3000 ns to 396 ns

– 50% more protons/bunch

• √s will increase from 1.8 to 2 TeV

• Limitation will be on p̄/bunch

• Plan on going to 132 ns bunch spacing later

The Main Injector construction is well underway

• Will be doing some machine studies in fall 1995

• Fixed target run in 1996

• Expect to start commissioning in late 1998

Future Collider Developments

Tevatron Run III has a goal of L = 10^{33} cm^{−2}s^{−1}

• Achieve much of this with increases in p̄ flux

• Would provide ∼ 10 fb−1 per year

• Competitive with LHC in early years of operation(?)

LHC will blow the socks off Tevatron once running

• Luminosity from 1033 to 1034 cm−2s−1

• Center of mass energy from 8 TeV to 14 TeV

• Enormous production cross section advantage

• Issue is only timescale (≥ 2004)

LHC programme combines both increased L and √s

• Both detectors (ATLAS and CMS) suited for top studies • Definitive top physics measurements will be done

• Huge source of t t events

• Form significant background to H → W + W − search

Limitations to Luminosity Increase

Significant limitations to increasing L

• Number of interactions/crossing becomes an issue • Can’t afford to build new detectors

• Have to adabiatically approach higher L

Run II CDF Detector Upgrades

CDF is being refurbished

• New electronics needed for shorter bunch spacing • Longer SVX detector with almost complete coverage

– Deadtimeless readout

• New calorimeter in Plug/Forward region – scintillator tile with fibre readout

• Enhanced μ coverage (fill gaps)

• Items under consideration – Central tracking system not likely to survive – Add scintillating fibre intermediate tracker – Particle ID system (TOF or RICH?)

Overall upgrade project is ∼ $40M effort

• Will be tough to get done in time

D0 / Upgrades for Run II

D0 / is metamorphising into a magnetic spectrometer

• Adding a solenoid and new central tracker

• Rebuilding front-end for 396 and 132 ns operation

• Adding SVX detector for b tagging

• Improvements in triggering

Collaboration eager to get this device on-line

• Makes it at least as “good” a detector as CDF

• Can compete in both the low and high pT physics programmes

Top Physics Impact

Biggest impact Run II will have is increased statistics

• ×5 − 10 in integrated luminosity

• More complete b tagging

Can look at the obvious physics measurements

• σt t

• Mtop

• Branching ratios and Vtb

• Search for new physics

Can make extrapolations based on Run I experience

Run II Top Yields

Can estimate what we should expect in next five years

• Use known yields for Run I

• W + ≥ 4 jet sample has S/N> ∼ 2

• Add in additional tagging capability

Channel  1 fb−1 10 pb−1
Tagged W + ≥ 4 jets  600  6000
Double tagged W + ≥ 4 jets  300  3000
Tagged Dilepton events  100  1000
Z+ ≥ 4 jet events  200  2000

Cross section measurement will scale with sensitivity

• Statistical uncertainty ∼ 5% range

• Most systematic uncertainties scale with sample size – integrated luminosity an exception – ∼ 3%

• Overall uncertainty ±5 (stat ) ± 7 (syst )% – Currently theoretical uncertainty is ∼ ±20%

How Well Can We Measure M_{top}?

Top mass statistical uncertainty will go like 1/√L

• Will go down to ∼ 2 GeV/c2 for Run II

• Reduce by using dileptons or multijet events?

Systematic uncertainty driven by jet reconstruction

• Best control is W → qq̄′ signal

• Also have Z + 1 and γ + 1 jet as one calibration

• Expect to reduce σjet to ∼ 5 GeV/c2 in Run I – Use W signal and higher statistics control samples

• Should scale as 1/√L – Experience of W mass instructive

– Should be able to get to ∼ 1.5 − 2.0 GeV/c2

Other uncertainties (6.4 GeV/c^{2}) should scale

• knowledge of background shape and b tagging bias

• These should reduce to ∼ 1.7 GeV/c2

• Gives total systemaatic uncertainty of 2.4 GeV/c2

δ(M_{top} ) ∼ 3 GeV/c^{2}

Run II M_{top} and M_{W} measurements will constrain SM

12

How About V_{tb}?

Limits on V_{tb} depend on V_{ts} and V_{td}

• Best to quote uncertainty on BR(t → W + b)

• Will remain dominated by statistical uncertainties – Biggest systematic uncertainty is tagging efficiency – Will have ∼ 107 B → lX events to measure this

Uncertainty from W + jets measurement is ∼ ±20%

• Should get to statistical uncertainty of ∼ ±3%

• Dileptons will measure BR to ∼ ±4%

• Should continue to scale with 1/√L

With earlier assumption about V_{ts} and V_{td}

• Set limit of Vtb > 0.25 at 95% CL

Searches for New Physics

The search for a new Z′ drive by statistics

• Can use “Top color” model to determine sensitivity • Will easily discover it if MZ′ < ∼ 1 TeV/c2

Single Top Production

Single top production is expected in SM

• Expected to be < 10% of t t cross section • Could be enhanced by non-SM physics

– For example, W ′ → tb̄

The W′ process is already interesting

W ′ → t b̄

|_{⟶}W + b

• Would see signal in W bb̄ invariant mass • Competitive with other W ′ searches

– Doesn’t have same S/N problem as W ′ → qq̄′ – Higher rate than leptonic modes

– Expect ∼ 60 fb of reconstructed tagged events for 500 GeV/c2 W ′

Top Physics at the LHC

Most of the detailed top studies for LHC are now dated

• Studies at Tevatron have taken centre stage

• Run II expectations approach LHC estimates

• But huge statistics allow different techniques – 6000 t t events/day at 1033 cm−2s−1

Can measure M_{top} using dilepton events

t → W + b → l′−ν̄l′X

|_{⟶}l+ νl

• Use l+ l′− invariant mass distribution

• Does not rely on jet reconstruction

• Insensitive to t pT

ATLAS expects to achieve accuracy of ∼ ±2 GeV/c^{2} in M_{top}

• Similar results from other studies (CMS and SSC)

• Will become systematics limited very quickly

Branching Ratios at LHC

ATLAS (CERN/LHCC/94-43) analysed top BR measurements

• Could get BR(t → W + b) measured to ∼ ±5% – Concerned about backgrounds from W bb̄ and gb → tW − – Didn’t consider making t t invariant mass cut

– Suspect this estimate is conservative

• Sensitive to BR(t → Zc) ∼ 10−5

Will also be sensitive to t → H^{+} b via H^{+} → τ^{+} ν_{τ} decays

• excess of τ jets in b-tagged W + jet events

• Can measure this to ∼ ±3%

• Sensitivity depends on tan β

Conclusions

The study of top at hadron colliders has just begun

• Have first results on 45 − 67 pb−1 – Mtop and t t kinematics

– Production cross section and branching ratios • Will have 50 tagged W ≥ 4 jet events for Run I

• Enormous increase in our understanding of – b-tagging for top physics

– Mass reconstruction using jets

– Backgrounds to t t production

Prospects for Top physics at Tevatron excellent

• Expect ×12 in sample size by the year 2000

• Run III might bring another ×10 in statistics

Action will shift to LHC in 2004-2005

• Becomes a “top factory” even in commissioning phase

• Study of top becomes essential for background control in the search for the Higgs and other new physics